I can't figure out the ones marked incorrect.

Burlington Clock Works manufactures fine, handcrafted clocks. The firm uses a job-order costing system, and manufacturing overhead is applied on the basis of direct-labor hours. Estimated manufacturing overhead for the year is $260,000. The firm employs 10 master clockmakers, who constitute the direct-labor force. Each of these employees is expected to work 2,000 hours during the year, which represents each employees practical capacity. The following events occurred during October. a. The firm purchased 3,500 board feet of mahogany veneer at $12 per board foot. b. Twenty brass counterweights were requisitioned for production. Each weight cost $24 c. Five gallons of glue were requisitioned for production. The glue cost $20 per gallon. Glue is treated as an indirect material d. Depreciation on the clockworks building for October was $9,000 e. A $600 utility bill was paid in cash f. Time cards showed the following usage of labor: Job number G60: 12 grandfathers clocks, 1,000 hours of direct labor Job number C81: 20 cuckoo clocks, 700 hours of direct labor The master clockmakers (direct-labor personnel) earn $18 per hour. g. The October property tax bill for $900 was received but has not yet been paid in cash h. The firm employs laborers who perform various tasks such as material handling and shop cleanup. Their wages for October amounted to $2,500 i. Job number G60, which was started in July, was finished in October. The total cost of the job was $14,400 j. Nine of the grandfathers clocks from job number G60 were sold in October for $1,500 each 2. Prepare journal entries to record the events described above. (If no entry is required for a transaction/event, select No journal entry required in the first account field.) Answer is complete but not entirely correct. No Transaction General Journal Debit Credit Raw-material inventory 42,00 Accounts payable 42,00 2 Work-in-process inventory Raw-material inventory 48 48 Manufacturing overhead Manufacturing-supplies inventory Manufacturing overhead 9,00 Accumulated depreciation: Building 9.00 Manufacturing overhead 60 Cash 60 6 Work-in-process inventory 18.00 Wages payable 18.00 f(2) Work-in-process inventory 14.00 Wages payable 14.00 8 Manufacturing overhead Property taxes payable 9 90 90 Manufacturing overhead 2,50 Wages payable 2,50 10 Finished-goods inventory 14,40 Work-in-process inventory 14,40 Wages payable Wages payable 12 j(2) Wages payable Wages payable
Answer #1
a overhead = 260000
b Number of hr =10*2000 20000
c=a/b Overhead rate per hr = 13
Journal entry Dr Cr
F(1) Work in progress inventory (G60) =18*1000 18000
Work in progress inventory (C81) =18*700 12600
Wage payable 30600
F(2) Work in progress inventory (G60) =13*1000 13000
Work in progress inventory (C81) =13*700 9100
Manufacturing overhead 22100
J(1)
Account receivable/Cash =9*1500 13500
Sales 13500
J(2) Cost of goods sold =14400*9/12 10800
Finished goods inventory 10800